# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 3. Gluing of Cauchy data and horizon exclusion
(25) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)=\bigg(\mbox{tr}\chi(-a,0)-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{C}{a^{2}}\bigg)$
(26) $-\bigg|\mbox{tr}\chi(-a,0)+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{C}{a^{2}}\bigg|.$
Now integrating the $\nabla_{4}$ equation one can control $|\underline{\hat{\chi}}|^{2}(u,\epsilon)$ in terms of $|\underline{\hat{\chi}}|^{2}(u,0)$. Recall the following $\nabla_{4}$ equation verified by $\underline{\hat{\chi}}$
@@
which, after integration and utilizing the estimates 15
@@
Therefore, $|u||\underline{\hat{\chi}}|^{2}(u,\epsilon)\approx|u||\underline{\hat{\chi}}|^{2}(u,0)$ up to negligible error term. Therefore, we observe the following
@@
(30) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=$ $\bigg(\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{C}{a^{2}}\bigg)$
@@
(31) $-\bigg|\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{C}{a^{2}}\bigg|.$
Subsequently, we would want
(32) $\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{C}{a^{2}}\approx 0$
which is compatible with the data choice. which would lead to the following two expressions
(33) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)=-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{C}{a^{2}}$

### v4

@@ 2.1. Yau [44] criterion for MOTS formation
(24) $2\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)$ $=\bigg(\mbox{tr}\chi(-a,0)-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}-\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg)$
$-\bigg|\mbox{tr}\chi(-a,0)+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg|.$
Integrating the $\nabla_{4}$ equation one can control $|\underline{\hat{\chi}}|^{2}(u,\epsilon)$ in terms of $|\underline{\hat{\chi}}|^{2}(u,0)$. Recall the following $\nabla_{4}$ equation verified by $\underline{\hat{\chi}}$
@@
which, after integration and utilizing the estimates (13)
@@
Therefore, $|u||\underline{\hat{\chi}}|^{2}(u,\epsilon)\approx|u||\underline{\hat{\chi}}|^{2}(u,0)$ up to a negligible error term at least $O(a^{-5/2})$ since each of the error terms decay much faster than $|u|^{-2}$. Therefore, we observe the following
@@
(28) $2\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=$ $=\bigg(\mbox{tr}\chi(-a,0)-\frac{4\epsilon}{a^{2}}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}-\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg)$
$-\bigg|\mbox{tr}\chi(-a,0)-\frac{4\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg|$
Now, our goal is to obtain a strictly positive lower bound for $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)$. But the second term can be potentially problematic. Note on the other hand that we prescribe regular data on $H_{u_{\infty}}$. Given $\underline{\hat{\chi}}$ on $\underline{H}_{0}$, the remaining Ricci coefficients are determined on $\underline{H}_{0}$. Therefore, one ought to integrate the $\nabla_{3}$ transport equation for $\mbox{tr}\chi$ to estimate $\mbox{tr}\chi(-a,0)$ using the data $\mbox{tr}\chi(u_{\infty},0)$. Subsequently, we would want
(29) $\mbox{tr}\chi(-a,0)+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}-\frac{4\epsilon}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}=O(a^{-\frac{5}{2}})$
which is compatible with the data choice. Now we ought to obtain $\mbox{tr}\chi(-a,0)$ in terms of its data $\mbox{tr}\chi(u_{\infty},0)$ since it is prescribed on $u=u_{\infty}$ hypersurface. Direct integration of the equation
$\nabla_{3}\mbox{tr}\chi+\frac{1}{2}tr\underline{\chi}\mbox{tr}\chi=2\underline{\omega}\mbox{tr}\chi+2\text{div}\eta+2|\eta|^{2}+2\rho-\hat{\chi}\cdot\hat{\underline{\chi}}$
yields
(30) $\mbox{tr}\chi(-a,0)=\frac{|u_{\infty}|\mbox{tr}\chi(u_{\infty},0)}{a}+2\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\bigg(\mbox{div }\eta+\rho-\frac{1}{2}\hat{\chi}\cdot\underline{\hat{\chi}}\bigg)(u^{{}^{\prime}},0)du^{{}^{\prime}}+O(a^{-\frac{5}{2}}).$
Importantly, $\mbox{div }\eta+\rho-\frac{1}{2}\hat{\chi}\cdot\underline{\hat{\chi}}$ is nothing but the negative of the mass aspect function $\mu$. Consequently
(31) $\mbox{tr}\chi(-a,0)=\frac{|u_{\infty}|\mbox{tr}\chi(u_{\infty},0)}{a}-\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}+O(a^{-\frac{5}{2}}).$
Now this forces the choice of $\mbox{tr}\chi(u_{\infty},0)$ since we want to keep the data $\underline{\hat{\chi}}$ on $\underline{H}_{0}$ free. Substituting $\mbox{tr}\chi(-a,0)$ from (30) into (29) yields
$\frac{|u_{\infty}|}{a}\mbox{tr}\chi(u_{\infty},0)-\frac{4\epsilon}{a^{2}}-\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}=\mathscr{E},$
where $\mathscr{E}=O(a^{-\frac{5}{2}})$. The idea here is that we prescribe the data on $H_{u_{\infty}}$ in terms of data on $H_{0}$ that is consistent with the null structure equations. This is because once we prescribe $\underline{\hat{\chi}}$ on $H_{0}$, the remaining Ricci coefficients are determined. This yields
@@
(32) $\bigg(\mbox{tr}\chi(u_{\infty},0)+\mbox{tr}\underline{\chi}(u_{\infty},0)\bigg)-\frac{4\epsilon}{|u_{\infty}|a^{2}}-\frac{2}{|u_{\infty}|}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu du^{{}^{\prime}}+\frac{a}{|u_{\infty}|}\bigg[\frac{|u_{\infty}|^{2}}{a^{2}}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})$
(33) $-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]=\frac{a}{|u_{\infty}|}\mathscr{E}.$
One aims for the following for $\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})$
(34) $\frac{|u_{\infty}|^{2}}{a^{2}}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})=\frac{1}{10a}\bigg[\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]$
since $\underline{\hat{\chi}}$ on $\underline{H}_{0}$ is free data and therefore
(35) $\bigg(\mbox{tr}\chi(u_{\infty},0)+\mbox{tr}\underline{\chi}(u_{\infty},0)\bigg)=\frac{4\epsilon}{|u_{\infty}|a^{2}}+\frac{2}{|u_{\infty}|}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{9}{10|u_{\infty}|}\bigg[\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
$+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]+\frac{a}{|u_{\infty}|}\mathscr{E}$
Now, in terms of the scaling (13)
(36) $\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}=O(a^{-\frac{1}{2}}),~\mathscr{E}=O(a^{-\frac{5}{2}}),~\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}=O(a^{-1}),$
which, in light of the estimates
(37) $|\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})|=O(|u_{\infty}|^{-2})~i.e.,~\bigg|\mbox{tr}\underline{\chi}(u_{\infty})+\frac{2}{|u_{\infty}|}\bigg|=O(|u_{\infty}|^{-2})$
yields the following asymptotics of $\mbox{tr}\chi$
(38) $\bigg|\mbox{tr}\chi(u_{\infty})-\frac{2}{|u_{\infty}|}\bigg|=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1})$
and a more precise value of $\mbox{tr}\chi(u_{\infty})$ would be
(39) $\mbox{tr}\chi(u_{\infty})=\frac{2}{|u_{\infty}|}+\frac{4\epsilon}{|u_{\infty}|a^{2}}+\frac{2}{|u_{\infty}|}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{9}{10|u_{\infty}|}\bigg[\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
$+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]$
up to an error term $O(a^{-\frac{3}{2}}|u_{\infty}|^{-1})$ and $a\gg 1$. Recall the following facts about the characteristic data on $\underline{H}_{0}$
  1. (1)
    The conformal class/metric on the corner sphere $S_{u_{\infty},0}$
  2. (2)
    The incoming shear $\underline{\hat{\chi}}$ prescribed on the null hypersurface $\underline{H}_{0}$.
  3. (3)
    A gauge normalization, for example: $\Omega$ fixed on $S_{u_{\infty},0}$, together with a choice of the shift vector $b$ along the initial null hypersurface $\underline{H}_{0}$.
There data are precisely provided on $\underline{H}_{0}$ and on the corner sphere $S_{u_{\infty},0}$. See chapter $2$ of [19]for the related concepts in their framework. This would lead to the following two expressions for the generalized mean curvature
(40) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)$ $=-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
(41) $+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}$